if your on a plane that went 480 miles... and for the first 240 miles you went 60 and the last 240 you did 50mph what was your average mph for the trip
i took the two mph and dived by 2 to get 55

Found 2 solutions by rakhileshnaidu, Alan3354:
Answer by rakhileshnaidu(13)
Answer by Alan3354(69443)

if your on a plane that went 480 miles... and for the first 240 miles you went 60 and the last 240 you did 50mph what was your average mph for the trip
i took the two mph and dived by 2 to get 55
-------------
Avg = 2*60*50/(60+50) = 6000/110
=~ 54.55 mi/hr
---------------
It's not the average of 50 & 60
---------
Find the total time:
t = 240/60 + 240/50 = 4 + 24/5
t = 44/5 hours
----
r = d/t = 480/(44/5) = 2400/44
= same answer
=================
It's not the avg of 60 & 50 because more time is spent moving at the lower speed.
If the 2 speeds were 60 & 30, the overall speed = 40, not 45.
This applies to round-trips, or trips of equal distance only.